# Diverse, not Short: A Length-Controlled Self-Learning Framework for Improving Response Diversity of Language Models

### v1

@@ 1 Introduction
This decrease in diversity has important consequences. When humans collaborate with aligned models, the content they produce tends to be less original and less varied (Doshi and Hauser, 2024; Padmakumar and He, 2023). At scale, this reduction in diversity can hinder creative ideation and increase output homogeneity (Anderson et al., 2024; Xu et al., 2024). Beyond creativity, reduced diversity of generated text has a direct impact on the continued improvement of LLMs. Recent studies have shown that repeatedly training models on their own aligned outputs can lead to a consistent decline in diversity, eventually resulting in model collapse (Shumailov et al., 2023; Guo et al., 2023; Seddik et al., 2024; Herel and Mikolov, 2024).
@@
    Length-controlled diversity evaluation: We introduce Diversity Decile, a new metric that adjusts for text length when evaluating diversity gains.
@@ 2 Related Work
Evaluation of diversity is challenging for two main reasons: length bias McCarthy and Jarvis (2010); Covington and McFall (2010); Mass (1972); Johnson et al. (2023), and inconsistent human preferences Evans et al. (2016); Chakrabarty et al. (2023, 2024); Gómez-Rodríguez and Williams (2023). Despite the challenges, many studies have highlighted the compromised diversity of synthetic text Shaib et al. (2024b, a); Salkar et al. (2022); Padmakumar and He (2023); Guo et al. (2023); Kirk et al. (2023); Doshi and Hauser (2024); Anderson et al. (2024). So, we present a method, Diverse-NS, to increase the response diversity and propose a metric, Diversity Decile, to measure diversity in a length-controlled way.
@@ 3 Preliminaries
Self-learning, also known as self-training, is a semi-supervised approach involving three main steps: data generation (pseudo-labeling), data filtering, and model learning (Lee et al., 2013; Amini et al., 2025). In our setup, data generation involves sampling text from a language model in response to story-writing prompts. This is followed by filtering, where we construct high-quality preference pairs—two continuations for the same prompt, with one preferred over the other. We refer to the preferred continuation as the “chosen” continuation (or response) and the other as the “rejected” continuation (or response). Using this preference dataset, we apply Direct Preference Optimization (DPO) (Rafailov et al., 2023) to train the model to favor the chosen responses.
@@ 5.2.2 Length-Adjusted Evaluation
While most diversity metrics exhibit bias toward shorter outputs, Johnson et al. (2023) shows that the DSI metric displays the opposite tendency-it favors longer responses. This is not an issue in tasks like DAT, where the output length is fixed at 10 words. But for open-ended tasks such as CWT, longer stories may receive disproportionately high DSI scores primarily due to their length, rather than genuine diversity. To address this issue, we introduce a novel evaluation metric: $\Delta$ Diversity Decile ($\Delta$ DD), which takes into account text length when assessing diversity.

### v2

@@ 1 Introduction
This decrease in diversity has important consequences. When humans collaborate with aligned models, the content they produce tends to be less original and less varied (Doshi and Hauser, 2024; Padmakumar and He, 2023). At scale, this reduction in diversity can hinder creative ideation and increase output homogeneity (Anderson et al., 2024; Xu et al., 2024). Beyond creativity, reduced diversity of generated text has a direct impact on the continued improvement of LLMs, with only limited benefits of reduced diversity (Deshpande et al., 2023; Muckatira et al., 2024). Recent studies have shown that repeatedly training models on their own aligned outputs can lead to a consistent decline in diversity, eventually resulting in model collapse (Shumailov et al., 2023; Guo et al., 2023; Seddik et al., 2024).
@@
    Length-controlled diversity evaluation: We introduce Diversity Decile, a diversity metric that adjusts for text length.
@@ 2 Related Work
Evaluation of diversity is challenging for a few reasons: length bias McCarthy and Jarvis (2010); Covington and McFall (2010); Mass (1972); Johnson et al. (2023), relative difficulty in achieving substantial agreement between humans (Chakrabarty et al., 2023, 2024; Gómez-Rodríguez and Williams, 2023), and inconsistent human preferences Evans et al. (2016). Despite the challenges, many studies have highlighted the compromised diversity of synthetic or human-LLM collaborative text Shaib et al. (2024b, a); Salkar et al. (2022); Padmakumar and He (2023); Guo et al. (2023); Kirk et al. (2023); Doshi and Hauser (2024); Anderson et al. (2024). So, we present a method, Diverse-NS, to increase the response diversity and propose a metric, Diversity Decile, to measure diversity in a length-controlled way.
@@ 3 Preliminaries
Self-learning, also known as self-training, is a semi-supervised approach involving three main steps: data generation (pseudo-labeling), data filtering, and model learning (Lee et al., 2013; Amini et al., 2025). In our setup, data generation involves sampling text from a language model in response to story-writing prompts. This is followed by filtering, where we construct high-quality preference pairs—two continuations for the same prompt, with one preferred over the other. We refer to the preferred continuation as the “chosen” and the other as the “rejected”. Using this preference dataset, we apply Direct Preference Optimization (DPO) (Rafailov et al., 2023) to train the model to favor the chosen responses.
@@ 5.2.2 Length-Adjusted Evaluation
While most diversity metrics exhibit bias toward shorter outputs, Johnson et al. (2023) shows that the DSI metric displays the opposite tendency-it favors longer responses. This is not an issue in tasks like DAT, where the output length is fixed at 10 words. But for open-ended tasks such as CWT, longer stories may receive disproportionately high DSI scores primarily due to their length, rather than genuine diversity. To address this issue, we introduce a novel evaluation metric: $\Delta$ Diversity Decile ($\Delta$ DD), which accounts for text length when assessing diversity.